# TransientVerse: A Comprehensive Real-Time Alert and Multi-Wavelength Analysis System for Transient Astronomical Events

### v1

@@ 1 Introduction
In order to facilitate the real-time distribution of these alert messages to other astronomers for follow-up observations using multi-wavelength and multi-messenger instruments, it is essential to establish a workflow that can process and analyze these alert messages, thereby enabling their distribution, filtering, and classification. Although there are already many excellent astronomical alert platforms such as Astronomer’s Telegram (Atel) [^1] https://astronomerstelegram.org/, Gamma-ray Coordinates Network (GCN) [^2] https://gcn.nasa.gov/notices, and Virtual Observatory Event (VOEvent) [^3] https://www.chime-frb.ca/voevents, these alerts are often published across multiple platforms, leading to incomplete information dissemination. Furthermore, some alert platforms do not use the machine-readable standard data formats developed by the International Virtual Observatory Alliance (IVOA), but instead release observations in human-readable formats, which poses challenges to batch processing and automation. To address the challenges of scattered, inconsistent, and non-retroactive alert messages, it is crucial to establish a platform that can integrate and push transient source alerts in real-time. This will significantly improve observational efficiency, enhance the comprehensive understanding of these transient phenomena, and promote near-real-time multi-wavelength follow-up observations of transient sources.
To address the growing challenges of observational data and alerts, we develop a real-time notification platform called “TransientVerse”. TransientVerse collects and integrates transient source alerts from multiple sources, such as Atel, GCN, and VOEvent, through an automated collection pipeline. By leveraging open-source large language models, the platform parses textual data into structured formats, which are promptly delivered to users via standardized templates, facilitating multi-wavelength and multi-messenger follow-up observations. Additionally, TransientVerse provides retrospective query, visualization, and literature research functionalities, allowing users to efficiently analyze the collected alerts.
@@ 3.1.2 Report Parsing
After the report acquisition process, the pipeline undergoes a standardized parsing procedure to extract relevant fields of information from the reports. This stage integrates Large Language Models (LLMs) with custom parsers to achieve high accuracy. The LLMs are fine-tuned with specific prompts to extract structured data, while custom parsers are responsible for handling particular file formats. Once the data is parsed, it is converted into a unified structure, ensuring that all extracted information follows a standardized format for efficient querying and analysis.
In this process, several core fields are selected for extraction, including Transient Name, Transient Type, Telescope, RA (Right Ascension), Dec (Declination). The evaluation of different models is based on a combination of accuracy, precision, and edit distance metrics. After extensive testing and comparison, GPT-4 Turbo (OpenAI, 2024) and ERNIE-Bot 4.0 (Baidu Research, 2024) were identified as the best-performing models for this task. These models demonstrated superior capabilities in parsing the reports and extracting the necessary data, providing a reliable solution for handling complex report formats and ensuring consistent data quality for downstream analysis.
@@ 3.2.1 Search for Transient Source Alert Messages
Figure 4: Structured alert data representation on TransientVerse. The Report Title column links to the original textual report, while the Source Name column provides detailed source information, including basic data, related alerts, and papers retrieved from ADS. Keywords such as RA, Dec, RA_hms, Dec_hms, Loc_err, Source Name, Report Subjects, Telescope, and Source Type are extracted from the textual messages using LLMs.
@@ 3.2.2 Display of Transient Source Alerts
  • Structured Data Display: Key information such as source name, coordinates, observation time, and telescope name, etc., is presented in a table (see Table 2). For human-readable texts from Atel and Circulars, fields like RA, Dec, RA_hms, Dec_hms, Loc_err, Source_Name, Report Subjects, Telescope, and Source_Type are extracted via open-source LLMs, though the extraction accuracy is not guaranteed. Users can filter, analyze, and download selected reports in CSV format. Clicking on the “source name” reveals basic information retrieved from the SIMBAD database, related reports, and papers from ADS. Clicking the “Report Title” opens the original full alert message.
@@ 3.2.6 Follow-up Observation Alert Feature for FRBs Based on VOEvent Messages
4 Summary and Outlook
The TransientVerse has been designed with the objective of becoming the world’s leading comprehensive platform for transient source alert management. It facilitates a range of services, including subscription management, retrospective queries, and rapid access to basic information, historical alert data, and related literature for sources referenced in the reports. In addition, the platform’s visualization tools and user-friendly interface support real-time queries and personalized customization, significantly optimizing the research workflow. The modular design of message sources ensures the platform’s high scalability, allowing for flexible integration with external services and data sources, thus providing strong support for multi-wavelength and multi-messenger astronomy research. TransientVerse platform presents several remarkable features and advantages as follows:
  • Integration and Structuring of Alert Messages: The platform integrates alert messages from various transient alert systems and utilizes open-source large models to convert textual alerts into structured data. This approach enhances data consistency and query efficiency, laying the foundation for further analysis and visualization. The modular design also facilitates the integration of new alert sources and telescope systems.
@@ 4 Summary and Outlook
  • Sky Map Visualization: The platform provides sky map-based visualization of transient sources, supporting multiple coordinate systems and projection modes. Users can overlay specific telescope visibility regions to quickly determine whether a target is within the observable range, thereby improving observation planning efficiency.
  • Source Investigation and Literature Support: By clicking on a transient source’s name, users can access detailed information, including historical alert records and related academic literature. The platform leverages the ADS system to perform precise searches for literature associated with the source, aiding researchers in obtaining reference materials efficiently.

### v2

@@ 1 Introduction
In order to facilitate the real-time distribution of these alert messages to other astronomers for follow-up observations using multi-wavelength and multi-messenger instruments, it is essential to establish a workflow that can process and analyze these alert messages, thereby enabling their distribution, filtering, and classification. Although there are already many excellent astronomical alert platforms such as Astronomer’s Telegram (Atel) [^1] https://astronomerstelegram.org/, Gamma-ray Coordinates Network (GCN) [^2] https://gcn.nasa.gov/, and Virtual Observatory Event (VOEvent) [^3] https://www.chime-frb.ca/voevents, these alerts are often published across multiple platforms, leading to incomplete information dissemination. Furthermore, some alert platforms do not use the machine-readable standard data formats developed by the International Virtual Observatory Alliance (IVOA), but instead release observations in human-readable formats, which poses challenges to batch processing and automation. To address the challenges of scattered, inconsistent, and non-retroactive alert messages, it is crucial to establish a platform that can integrate and push transient source alerts in real-time. This will significantly improve observational efficiency, enhance the comprehensive understanding of these transient phenomena, and promote near-real-time multi-wavelength follow-up observations of transient sources.
To address the growing challenges of observational data and alerts, we develop a real-time notification platform called “TransientVerse” [^4] https://transientverse.zero2x.org/. TransientVerse collects and integrates transient source alerts from multiple sources, such as Atel, GCN, and VOEvent, through an automated collection pipeline. By leveraging open-source large language models, the platform parses textual data into structured formats, which are promptly delivered to users via standardized templates, facilitating multi-wavelength and multi-messenger follow-up observations. Additionally, TransientVerse provides retrospective query, visualization, and literature research functionalities, allowing users to efficiently analyze the collected alerts.
@@ 3.1.2 Report Parsing
A standardized parsing procedure follows the report acquisition process to extract relevant information from the reports. This stage integrates Large Language Models (LLMs) with custom parsers to ensure high accuracy. The LLMs are fine-tuned using carefully designed prompts, while custom parsers handle specific file formats. Once the data is parsed, it is converted into a unified structure, ensuring that all extracted information follows a standardized format for efficient querying and analysis.
To optimize the performance of the LLMs in this text parsing task, we meticulously design prompts that include key components: role specification, task description, main objectives (with particular emphasis on fields like telescope names), and a well-crafted one-shot learning example. The role specification clarifies the model’s function, guiding it to identify and extract structured data from scientific reports. The task description explicitly outlines the objectives, such as extracting transient names, telescope identifiers, and astronomical coordinates (RA and Dec). Special attention is given to handling variations in the format of critical fields like telescope names, ensuring the model can correctly interpret and process diverse naming conventions.
A parsing example is included in the prompt to provide a clear demonstration of how the model should interpret the text. This one-shot learning approach allows the LLM to learn from a single example and apply that knowledge across the entire report. The input-output format is also clearly defined, ensuring that the model’s responses are structured and consistent.
In this process, several core fields are selected for extraction, including Transient Name, Transient Type, Telescope, RA (Right Ascension), and Dec (Declination). The effectiveness of different models is evaluated based on accuracy, precision, and edit distance metrics. Through extensive testing and comparison, GPT-4 Turbo (OpenAI, 2024) and ERNIE-Bot 4.0 (Baidu Research, 2024) were identified as the top-performing models for this task. These models demonstrated exceptional capabilities in parsing complex report formats and extracting the necessary data, offering a reliable solution that ensures consistent data quality for downstream analysis.
By combining fine-tuned LLM prompts with custom parsing techniques and a unified data structure, this approach significantly enhances the accuracy and efficiency of extracting structured data, facilitating seamless integration into subsequent analysis stages.
@@ 3.2.1 Search for Transient Source Alert Messages
Figure 4: Structured alert data representation on TransientVerse. The Report Title column links to the original textual report, while the Source Name column provides detailed source information, including basic data, related alerts, and papers retrieved from Astrophysics Data System (ADS) [^7] https://ui.adsabs.harvard.edu/. Keywords such as RA, Dec, RA_hms, Dec_hms, Loc_err, Source Name, Report Subjects, Telescope, and Source Type are extracted from the textual messages using LLMs.
@@ 3.2.2 Display of Transient Source Alerts
  • Structured Data Display: Key information such as source name, coordinates, observation time, and telescope name, etc., is presented in a table (see Table 2). For human-readable texts from Atel and Circulars, fields like RA, Dec, RA_hms, Dec_hms, Loc_err, Source_Name, Report Subjects, Telescope, and Source_Type are extracted via open-source LLMs, though the extraction accuracy is not guaranteed. Users can filter, analyze, and download selected reports in CSV format. Clicking on the “source name” reveals basic information retrieved from the SIMBAD [^8] https://simbad.u-strasbg.fr/simbad/sim-fbasic database, related reports, and papers from ADS. Clicking the “Report Title” opens the original full alert message.
@@ 3.2.6 Follow-up Observation Alert Feature for FRBs Based on VOEvent Messages
4 Discussion and Conclusion
The TransientVerse has been designed with the objective of becoming the world’s leading comprehensive platform for transient source alert management. It facilitates a range of services, including subscription management, retrospective queries, and rapid access to basic information, historical alert data, and related literature for sources referenced in the reports. In addition, the platform’s visualization tools and user-friendly interface support real-time queries and personalized customization, significantly optimizing the research workflow. The modular design of message sources ensures the platform’s high scalability, allowing for flexible integration with external services and data sources, thus providing strong support for multi-wavelength and multi-messenger astronomy research.
In the field of transient source astronomy, another platform similar to TransientVerse is Astro-COLIBRI Reichherzer et al. (2021) [^10] https://astro-colibri.org/. Although Astro-COLIBRI shares many features with TransientVerse, its design philosophy and application focus seem to differ in several ways. Astro-COLIBRI appears to place a greater emphasis on the detailed analysis of individual alert events. The event detail pages feature a discussion area where astronomers can exchange insights and share opinions on specific astronomical events, potentially fostering academic discourse. Additionally, the platform offers a mobile app, which further extends its accessibility and usability. In contrast, TransientVerse emphasizes the in-depth exploration of astronomical sources associated with alert events. By leveraging LLMs and custom parsers, the platform extracts source names from alert messages, retrieves essential source information from SIMBAD, searches related alert data within its database, and gathers relevant literature from ADS. This facilitates a rapid and comprehensive understanding of the sources involved in each event. Moreover, the platform features an observational assistant tool, which aids astronomers in determining the necessity of follow-up observations, thereby enabling multi-wavelength and multi-messenger collaborative efforts that drive progress in astrophysical discoveries.
TransientVerse platform presents several remarkable features and advantages as follows:
  • Integration and Structuring of Alert Messages: The platform integrates alert messages from various transient alert systems and employs open-source LLMs and custom parsers to convert textual alerts into structured data. This improves data consistency and query efficiency, laying the foundation for further analysis and visualization. Its modular design also supports the addition of new alert sources and telescope systems.
@@ 4 Discussion and Conclusion
  • Sky Map Visualization: The platform provides a sky map-based visualization of transient sources, supporting multiple coordinate systems and projection modes. Users can overlay specific telescope visibility regions to quickly determine whether a target is within the observable range, thereby improving observation planning efficiency.
  • Source Investigation and Literature Support: By clicking on a transient source’s name, users can access detailed information, including historical alert records and related academic literature. The platform leverages the ADS system to perform precise searches for the literature associated with the source, helping researchers to obtain reference materials efficiently.